How much of an 20% acid solution and how much of a 60% acid solution should bw mixed together to make 100 liters of a 35% acid solution?
Answer by ankor@dixie-net.com(22740)

How much of an 20% acid solution and how much of a 60% acid solution should be mixed together to make 100 liters of a 35% acid solution?
:
let x = amt of 60% solution
the total result is 100 l, therefore:
100-x = amt of 20% solution
:
decimal equiv of percent
.60x + .20(100-x) = .35(100)
.6x + 20 - .2x = 35
.6x - .2x = 35 - 20
.4x = 15
x = 15/.4
x = 37.5 liters of 60% solution required
and
100 - 37.5 = 62.5 liters of 20% solution
:
;
Check out solutions
.6(37.5) + .2(62.5) = 35
22.5 + 12.5 = 35